Errors in PostScript^{1}

Errors

dictfull  no more room in dictionary
dictstackoverflow  too many begins
dictstackunderflow too many ends
execstackoverflow exec nesting too deep
handleerror  called to report error information
interrupt  external interrupt request
invalidaccess  attempt to violate access atribute
invalidexit  exit not in loop
invalidfileaccess  unacceptable access string
invalidfont  invalid font name or dict
invalidrestore  improper restore
ioerror  input/output error occurred
limitcheck  implementation limit occurred
nocurrentpoint  current point is undefined
rangecheck  operand out of bounds
stackoverflow  operand stack overflow
stackunderflow  operand stack underflow
syntaxerror  syntax error in PostScript program text
timeout  time limit exceeded
typecheck  operand of wrong type
undefined  name not known
undefinedfilename file not found
undefinedresult  over/underflow or meaningless result
unmatched mark  expected mark not on stack
unregistered  internal error
VMERROR  VM exhausted

1Copied from the PostScript Language Reference Manual, Adobe Systems Incorporated.